# Native LLM and MLLM Inference at Scale on Apple Silicon

### v1

@@ Abstract.
The growing adoption of Apple Silicon for machine learning development has created demand for efficient inference solutions that leverage its unique unified memory architecture. However, existing tools either lack native optimization (PyTorch MPS) or focus solely on text models (llama.cpp), leaving multimodal workloads underserved.
We present vllm-mlx, a framework for efficient LLM and MLLM inference on Apple Silicon built natively on MLX. For text models, we achieve 21% to 87% higher throughput than llama.cpp across models ranging from Qwen3-0.6B to Nemotron-30B, while providing continuous batching that scales to 4.3x aggregate throughput at 16 concurrent requests. For multimodal models, we introduce content-based prefix caching that eliminates redundant vision encoding by identifying identical images through content hashing, regardless of input format.
Our evaluation on Apple M4 Max demonstrates throughput of up to 525 tokens per second on text models and 28x speedup on repeated image queries, reducing multimodal latency from 21.7 seconds to under 1 second. Video analysis with up to 64 frames achieves 24.7x cache speedup. We release our implementation as open source to support efficient inference on consumer Apple hardware.
@@ 1. Introduction
However, existing inference solutions for Apple Silicon face significant limitations. PyTorch’s MPS backend adapts CUDA-style operations to Metal but lacks native optimization for the unified memory model. llama.cpp provides excellent performance for text models but does not support vision-language models. vLLM-metal (vLLM Project, 2024), developed concurrently with our work, provides continuous batching but lacks multimodal support and vision caching. This fragmented landscape leaves developers without a unified solution for both text and multimodal inference on Apple Silicon.
@@ 2.3. LLM Inference on Apple Silicon
The current landscape for LLM inference on Apple Silicon includes several options. llama.cpp (Gerganov, 2023) provides highly optimized inference through hand-tuned Metal kernels and GGUF quantization, achieving excellent single-stream throughput. However, it lacks continuous batching for serving multiple concurrent requests. PyTorch with MPS backend offers broad model compatibility but suboptimal performance due to its CUDA-centric design. vLLM-metal (vLLM Project, 2024), developed concurrently with our work, provides an MLX-based plugin for vLLM with continuous batching support.
@@ 4.4. Discussion
Our framework currently supports only Apple Silicon, limiting deployment to macOS environments. Model support depends on mlx-lm and mlx-vlm availability. Future work includes: speculative decoding for improved single-stream latency, integration with vLLM’s upstream scheduler for feature parity, and support for additional multimodal architectures including audio models.
@@ 6. Conclusion
Future Work.
Next steps include extending caching to audio for speech-enabled MLLMs, exploring distributed inference across multiple Apple Silicon devices, and energy profiling for battery-powered deployments.

### v2

@@ Abstract.
The growing adoption of Apple Silicon for machine learning development has created demand for efficient inference solutions that leverage its unique unified memory architecture. However, existing tools either lack native optimization (PyTorch MPS) or focus solely on text models, leaving multimodal workloads underserved. We present vllm-mlx, a framework for efficient LLM and MLLM inference on Apple Silicon built natively on MLX. For text models, we achieve 21% to 87% higher throughput than llama.cpp across models ranging from Qwen3-0.6B to Nemotron-30B, while providing continuous batching that scales to 4.3x aggregate throughput at 16 concurrent requests. For multimodal models, we introduce content-based prefix caching that eliminates redundant vision encoding by identifying identical images through content hashing, regardless of input format. Our evaluation on Apple M4 Max demonstrates throughput of up to 525 tokens per second on text models and 28x speedup on repeated image queries, reducing multimodal latency from 21.7 seconds to under 1 second. Video analysis with up to 64 frames achieves 24.7x cache speedup. We release our implementation as open source to support efficient inference on consumer Apple hardware.
@@ 1. Introduction
However, existing inference solutions for Apple Silicon face significant limitations. PyTorch’s MPS backend adapts CUDA-style operations to Metal but lacks native optimization for the unified memory model. llama.cpp provides excellent performance for text models but does not support vision-language models. vLLM-metal (vLLM Project, 2024), the official vLLM backend for Apple Silicon, provides continuous batching but lacks multimodal support and vision caching. This fragmented landscape leaves developers without a unified solution for both text and multimodal inference on Apple Silicon.
@@ 2.3. LLM Inference on Apple Silicon
The current landscape for LLM inference on Apple Silicon includes several options. llama.cpp (Gerganov, 2023) provides highly optimized inference through hand-tuned Metal kernels and GGUF quantization, achieving excellent single-stream throughput. However, it lacks continuous batching for serving multiple concurrent requests. PyTorch with MPS backend offers broad model compatibility but suboptimal performance due to its CUDA-centric design. vLLM-metal (vLLM Project, 2024), the official vLLM backend for Apple Silicon, provides an MLX-based plugin with continuous batching support. Our work differs by adding native multimodal inference with vision-language models and content-based prefix caching that eliminates redundant vision encoding across conversation turns.
@@ 4.4. Discussion
Our framework currently supports only Apple Silicon, limiting deployment to macOS environments. Model support depends on MLX community. Future directions include speculative decoding for improved single-stream latency, distributed inference across multiple Apple Silicon devices via network-connected Mac clusters, and energy profiling for battery-powered deployments. We also plan to extend our caching approach to audio modalities for speech-enabled multimodal applications, and explore tensor parallelism across the GPU cores available in higher-end Apple Silicon configurations.